OpenStudy (anonymous):

Check my answer please :) The sum of the measures of the interior angles is (n - 2) 180 = (10 - 2) 180 = 8 x 180 = 1440 1440 <- my final answer @jim_thompson5910 @shinebrightlikeadimon

OpenStudy (anonymous):

yes that looks correct to me! good job btw are you online student

jimthompson5910 (jim_thompson5910):

I would show how you got n = 10 E = 360/n ... E is the exterior angle n*E = 360 n = 360/E n = 360/36 n = 10

jimthompson5910 (jim_thompson5910):

other than that, it looks good

OpenStudy (anonymous):

i agree with @jim_thompson5910

OpenStudy (anonymous):

oh right, so my final answer is correct?

OpenStudy (anonymous):

yes

OpenStudy (jenniferjuice):

@shinebrightlikeadimon COULD U PLEASE HELP ME ?

OpenStudy (jenniferjuice):

sorry forgot caps was on oops ...

OpenStudy (anonymous):

yes and np tag me in the question plz @jenniferjuice

OpenStudy (anonymous):

^thank you :)